# Distributed Certifiably Correct Range-Aided SLAM

### v1

@@ I Introduction
Figure 1: (a) Graphical representation of the RA-SLAM problem in the distributed setting. In DCORA, synchronized agents (e.g., $\mathcal{A}_{\alpha},\mathcal{A}_{\beta},\mathcal{A}_{\gamma}$) jointly estimate their state in a common frame by performing local computation steps and exchanging information over the distributed communication graph shown in (b). States, including poses (circles) and landmarks (stars), are interconnected via relative pose (black lines) and range measurements (blue lines). To facilitate efficient communication and computation, states are labeled private (black) or public (red), while measurements exist as inter-agent loop closures (dotted lines), intra-agent loop closures (dash-dotted lines), or odometry (solid lines) as discussed in Section III-C.
Specifically, DCORA applies distributed Riemannian optimization techniques [8] to solve the (convex) semidefinite programming (SDP) relaxation of a maximum likelihood (ML) formulation of RA-SLAM [9]. In turn, solving the SDP allows DCORA to recover globally optimal RA-SLAM solutions within practically relevant noise regimes. We solve the SDP via the Riemannian Staircase method [10, 11], for which we use the Riemannian block coordinate descent (RBCD) algorithm [8] to identify first-order critical points (i.e., points where the Riemannian gradient of the objective function vanishes). We leverage established Karush-Kuhn-Tucker (KKT)-based conditions (which require a certain dual certificate matrix, when evaluated at the global minima, to be positive semidefinite) of the RA-SLAM problem [12, 9] to verify global optimality, and apply distributed saddle escape techniques to descend from suboptimal critical points (if necessary). We benchmark DCORA against a state-of-the-art centralized certifiably correct RA-SLAM algorithm on real-world datasets, demonstrating comparable absolute trajectory error (ATE) performance. Furthermore, we perform parameter sweeps across synthetic datasets to observe how common parameters (namely the number of agents, the number of landmarks, and range measurement density) in RA-SLAM problems affect DCORA’s performance in terms of convergence rate, solution precision, and SDP tightness. These contributions are summarized as follows:
@@
    The first distributed RA-SLAM algorithm capable of producing certifiably globally optimal estimates;
@@
    A benchmark comparison of our algorithm against a state-of-the-art centralized certifiably correct RA-SLAM algorithm on real-world datasets;
@@
    A parametric study on DCORA’s convergence rate, solution quality, and SDP tightness using synthetic datasets;
@@ II-A Certifiably Correct Perception via the Riemannian Staircase
DCORA depends on a formulation of the RA-SLAM problem that allows for an SDP relaxation to be constructed and subsequently solved through a series of rank-restricted SDPs of considerably lower dimension than the original SDP. Furthermore, these specific rank-restricted SDPs are equivalent to optimization over the Stiefel and Oblique manifolds, leading DCORA to instead solve a series of (smooth) manifold optimization problems. This approach – solving an SDP via a series of lower-dimensional manifold optimization problems – is known as the Riemannian Staircase method [10, 11].
@@ II-C Distributed RA-SLAM Algorithms
Distributed SLAM algorithms have been developed to solve specific instances of the RA-SLAM problem. Such algorithms have been developed to fuse range measurements between robots [23, 24, 25, 26, 27] and into compound relative pose estimates between robots [28, 29, 30, 31, 32]; however, none provide a unified approach across all state and measurement combinations that arise in RA-SLAM. Moreover, these methods provide no guarantees (a priori or a posteriori) on the quality of the returned estimate. We note that several distributed back-end solvers [33, 34, 35, 36, 37, 38, 39] are capable of solving the RA-SLAM problem; however, they too lack formal guarantees on solution quality. Namely, these methods rely on local search techniques (e.g., Gauss-Newton [40]) and are thus susceptible to significantly suboptimal critical points. Even worse, certain local search techniques (e.g., Gaussian belief propagation [41]) lack convergence guarantees. The limitations of existing distributed RA-SLAM algorithms and distributed back-end solvers thus necessitate a new approach for safety-critical navigation.
@@ III-A RA-SLAM Problem Formulation and the Riemannian Staircase
To solve the RA-SLAM problem via the Riemannian Staircase, we begin with the following ML formulation of RA-SLAM:
@@
Given relative pose measurements $\{\tilde{R}_{ij},\tilde{t}_{ij}\}_{(i,j)\in\mathcal{E}_{p}}$ and point-to-point distance measurements $\{\tilde{r}_{ij}\in\mathbb{R}\}_{(i,j)\in\mathcal{E}_{r}}$, solve:
@@
Without loss of generality, 1 does not distinguish landmark variables from poses (as discussed in [9, Appendix E]) to simplify mathematical presentation. To recover globally optimal solutions to 1, [9] developed a (non-convex) quadratically constrained quadratic programming (QCQP) relaxation of 1, which enabled a (convex) semidefinite relaxation to be constructed via Shor’s relaxation [42]. The QCQP relaxation is as follows:
@@
Given the measurements of 1, find the variables, $\{R_{i},t_{i}\}_{i=1}^{n}\cup\{r_{ij}\}_{(i,j)\in\mathcal{E}_{r}}$ , which solve:
@@
2 relaxes the special orthogonal constraint $R_{i}\in\operatorname{{\operatorname{SO}}(d)}$ to an orthogonality constraint, $R_{i}\in\operatorname{{\operatorname{O}}(d)}$, and introduces auxiliary unit-norm vector variables, $r_{ij}\in\mathbb{R}^{d}$, as a reformulation of the range cost terms. From 2, an SDP relaxation can be constructed via Shor’s relaxation [42]:
@@
Find $Z\in\mathbb{R}^{k\times k}$ that solves:
@@
The exact steps to form this relaxation are outlined in [9, Appendix D]. Key to our approach is recognition that 3 should often admit low-rank solutions (i.e., solutions $Z^{*}$ for which $\operatorname{rank}Z^{*}\ll k$). Indeed, this low-rank behavior has been found in practice [9]. Thus, instead of solving 3 directly via standard SDP solvers, one may instead leverage the low-rank structure by optimizing over a low-dimensional factorization $X^{\top}X=Z,X\in\mathbb{R}^{p\times k}$ for user-selectable parameter $p\ll k$. This technique, known as the Burer-Monteiro method [43, 44, 45], arrives at the following problem:
@@
in which the factorization implicitly constrains the rank of the solution to be less than or equal to $p$. For our specific problem, $X$ can be structured as,
@@
For higher problem dimensions ($p>d$), variables (e.g., $R_{1}$) in (5) can be thought of as higher dimensional generalizations of the original variables of the QCQP (2); 4 thus represents a higher dimensional relaxation of 2. In fact, analysis of the constraints of 4 reveal that the problem can be equivalently posed as Riemannian manifold optimization [9]:
@@
Given the measurements of Problem 1, find the variables, $\{R_{i},t_{i}\}_{i=1}^{n}\cup\{r_{ij}\}_{(i,j)\in\mathcal{E}_{r}}$ , which solve:
@@
This manifold optimization representation (5) directly generalizes the original QCQP (2), as the Stiefel manifold $\operatorname{St}(p,d)$ generalizes the Orthogonal group $\operatorname{{\operatorname{O}}(d)}$ and the auxiliary variables $r_{ij}$ are still constrained to be unit-norm but can be of arbitrary dimension $p$. 5 is directly optimized in the DCORA algorithm. For notational convenience, we refer to the product manifold defined by 5’s variables as $\operatorname{\operatorname{\mathcal{M}}_{\text{RA-SLAM}}}(p,d)\triangleq\operatorname{St}(p,d)^{n}\times(S^{p-1})^{l}\times\mathbb{R}^{p\times n}\subset\mathbb{R}^{p\times k}$.
@@ III-B DCORA Algorithm
1:  Obtain initial point $X_{\text{init}}\in\operatorname{\operatorname{\mathcal{M}}_{\text{RA-SLAM}}}(p,d)$ for $p=d$ via distributed initialization.
@@
4:     Compute first-order critical point of 4 [8, Algorithm 3]: $X^{\ast}\leftarrow\text{RBCD}(X)$.
@@
6:     Construct corresponding dual certificate matrix [9, Theorem 1]: $S(X^{\ast})\triangleq Q+\sum_{i=1}^{m}\lambda_{i}A_{i}$, where $\lambda\in\mathbb{R}^{m}\triangleq\left[\lambda_{1},\ldots,\lambda_{m}\right]$ are the Lagrange multipliers at KKT point $X^{\ast}$.
7:     Compute minimum eigenpair [8, Algorithm 6]: $(\lambda,v)\leftarrow\text{MinEig}(S(X^{\ast}))$.
@@
11:        Construct second-order descent direction: $\dot{X}_{+}\triangleq[0\ v^{\top}]^{\top}$ where $v\in\mathbb{R}^{k}$ is the eigenvector corresponding to the smallest eigenvalue of $S(X^{\ast})$ which satisfies $v^{\top}S(X^{\ast})v<0$.
12:        Descend from $X^{\ast}$ [8, Algorithm 7]: $X\leftarrow\text{EscapeSaddle}(X^{\ast},\dot{X}_{+})$.
@@
15:  Recover the optimal value of the SDP relaxation: $f^{\ast}_{\text{SDP}}=\langle Q,{X^{\ast}}^{\top}X^{\ast}\rangle_{F}$, where $\langle\cdot,\cdot\rangle_{F}$ is the Frobenius inner product.
@@
An initial point $X_{\text{init}}\in\operatorname{\operatorname{\mathcal{M}}_{\text{RA-SLAM}}}(p=d,d)$ is obtained via distributed initialization. This initial point is lifted to rank $p_{0}$: $X=Y_{\text{rand}}X_{\text{init}}$, where $Y_{\text{rand}}\in\operatorname{St}(p_{0},d)$ is a random Stiefel element sampled from the initial rank $p_{0}\geq d$ of the Riemannian staircase.
From this lifted initial point, the Riemannian Staircase methodology is applied. First, 5 is optimized over via RBCD. Once RBCD has returned a first-order critical point $X^{\ast}$, $X^{\ast}$ is lifted to the next rank $p+1$ in the Riemannian staircase, where a distributed verification step either certifies $X^{\ast}$ as globally optimal or constructs a second-order descent direction to descend from $X^{\ast}$. If $X^{*}$ is not globally optimal, then saddle point escape is performed via line search along the computed descent direction and RBCD is used to optimize from the resulting point. The distributed Riemannian Staircase algorithm continues until a certified globally optimal value $f^{\ast}_{\text{SDP}}$ of 3 is recovered, which is a lower bound on 1’s optimal value. If the SDP is tight, as it often is in practice [9], then the solution is globally optimal for 1. Completing the DCORA algorithm, $X^{\ast}$ is projected to the feasible set of 1 via distributed rounding to recover pose and landmark estimates $T\in\operatorname{{\operatorname{SE}}(d)}^{n}\triangleq\{R_{i}\in\operatorname{{\operatorname{SO}}(d)},t_{i}\in\mathbb{R}^{d}\}_{i=1}^{n}$.
@@ III-C1 Variable Partitioning
A key notion in DCORA is that of variable ownership. Each decision variable in 5 is owned by a single agent and variables are partitioned by agent ownership to satisfy RBCD’s input requirement. Specifically, RBCD requires a block partitioning of $X$, where a block $X_{b}\subseteq X$ is associated with the agent responsible for estimating those states. In DC2-PGO, agents naturally take ownership of pose variables describing their trajectory. With the introduction of external landmarks and range measurements, variable partitioning becomes less obvious for the RA-SLAM problem. As such, we maintain DC2-PGO’s pose ownership rule in DCORA and assume agents own landmarks according to a consensus step [47] and auxiliary unit-norm vector variables corresponding to range measurements collected by their onboard range sensors.
@@ III-C2 Variable Labeling
In an agent’s local search step, RBCD requires agents to maintain certain local copies of variables belonging to other agents. To facilitate efficient sharing of these copies, poses and landmarks are labeled public or private according to the following definition:
@@ IV Experiments
Experiments conducted to evaluate DCORA are two-fold. First, we benchmark DCORA against its centralized counterpart, CORA, on public real-world multi-agent RA-SLAM datasets. Comparison is limited to CORA as (to our knowledge) this is the only other certifiably correct algorithm that can solve the general RA-SLAM problem. We refer the interested reader to [9] for an in-depth comparison of CORA against the GTSAM Levenberg–Marquardt optimizer [48], a popular local search back-end solver, to motivate our comparison. Second, a series of multi-agent RA-SLAM datasets are generated by modifying public synthetic PGO datasets. These datasets are modified to assign pose variables to multiple agents, include landmarks, and establish inter- and intra-agent range measurement loop closures with controlled density. Parameter sweeps are performed based on (i) the number of agents, (ii) the number of landmarks, and (iii) range measurement density to study the RA-SLAM problem structure. DCORA’s implementation for evaluation precedes the discussion of these experiments.

### v2

@@ I Introduction
Figure 1: (a) Graphical representation of the RA-SLAM problem in the distributed setting. In DCORA, synchronized agents (e.g., $\mathcal{A}_{\alpha},\mathcal{A}_{\beta},\mathcal{A}_{\gamma}$) jointly estimate their state in a common frame by performing local computation steps and exchanging information over the distributed communication graph shown in (b). States, including poses (circles) and landmarks (stars), are interconnected via relative pose (black lines) and range (blue lines) measurements. To facilitate efficient communication and computation, states are labeled private (black) or public (red), while measurements exist as inter-agent loop closures (dotted lines), intra-agent loop closures (dash-dotted lines), or odometry (solid lines) as discussed in Section III-C.
Specifically, DCORA applies distributed Riemannian optimization techniques [8] to solve the (convex) semidefinite programming (SDP) relaxation of a maximum likelihood (ML) formulation of RA-SLAM [9]. In turn, solving the SDP allows DCORA to recover globally optimal RA-SLAM solutions within practically relevant noise regimes. We solve the SDP via the Riemannian Staircase method [10, 11], for which we use the Riemannian block coordinate descent (RBCD) algorithm [8] to identify first-order critical points (i.e., points where the Riemannian gradient of the objective function vanishes). We leverage established Karush-Kuhn-Tucker (KKT)-based conditions (which require a certain dual certificate matrix, when evaluated at the global minima, to be positive semidefinite) of the RA-SLAM problem [12, 9] to verify global optimality, and apply distributed saddle escape techniques to descend from suboptimal critical points (if necessary). We benchmark DCORA against a state-of-the-art centralized certifiably correct RA-SLAM algorithm on real-world datasets, demonstrating comparable absolute trajectory error (ATE) performance. Furthermore, we perform parameter sweeps across synthetic datasets to observe how common parameters (namely, the number of agents, the number of landmarks, and range measurement density) in RA-SLAM problems affect DCORA’s performance in terms of convergence rate, solution precision, and SDP tightness. These contributions are summarized as follows:
@@
    The first distributed RA-SLAM algorithm capable of producing certifiably globally optimal estimates.
@@
    A benchmark comparison of our algorithm against a state-of-the-art centralized certifiably correct RA-SLAM algorithm on real-world datasets.
@@
    A parametric study on DCORA’s convergence rate, solution quality, and SDP tightness using synthetic datasets.
@@ II-A Certifiably Correct Perception via the Riemannian Staircase
DCORA depends on a formulation of the RA-SLAM problem that allows an SDP relaxation to be constructed and subsequently solved through a series of rank-restricted SDPs of considerably lower dimension than the original SDP. Furthermore, these specific rank-restricted SDPs are equivalent to optimization over the Stiefel and Oblique manifolds, leading DCORA to instead solve a series of (smooth) manifold optimization problems. This approach – solving an SDP via a series of lower-dimensional manifold optimization problems – is known as the Riemannian Staircase method [10, 11].
@@ II-C Distributed RA-SLAM Algorithms
Distributed SLAM algorithms have been developed to solve specific instances of the RA-SLAM problem. Such algorithms have been developed to fuse range measurements between robots [23, 24, 25, 26, 27] and into compound relative pose estimates between robots [28, 29, 30, 31, 32]; however, none provide a unified approach across all state and measurement combinations that arise in RA-SLAM. Moreover, these methods provide no guarantees (a priori or a posteriori) on the quality of the returned estimate. We note that several distributed back-end solvers [33, 34, 35, 36, 37, 38, 39] are capable of solving the RA-SLAM problem; however, they too lack formal guarantees on solution quality. That is, these methods rely on local search techniques (e.g., Gauss-Newton [40]) and are thus susceptible to significantly suboptimal critical points. Even worse, certain local search techniques (e.g., Gaussian belief propagation [41]) lack convergence guarantees. The limitations of existing distributed RA-SLAM algorithms and distributed back-end solvers thus necessitate a new approach for safety-critical navigation.
@@ III-A RA-SLAM Problem Formulation and the Riemannian Staircase
To solve the RA-SLAM problem via the Riemannian Staircase, we begin with the following ML formulation of RA-SLAM
@@
Given relative pose measurements $\{\tilde{R}_{ij},\tilde{t}_{ij}\}_{(i,j)\in\mathcal{E}_{p}}$ and point-to-point distance measurements $\{\tilde{r}_{ij}\in\mathbb{R}\}_{(i,j)\in\mathcal{E}_{r}}$, solve
@@
Without loss of generality, 1 does not distinguish landmark variables from poses (as discussed in [9, Appendix E]) to simplify mathematical presentation. To recover globally optimal solutions to 1, [9] developed a (non-convex) quadratically constrained quadratic programming (QCQP) relaxation of 1, which enabled a (convex) semidefinite relaxation to be constructed via Shor’s relaxation [42]. The QCQP relaxation is as follows
@@
Given the measurements of 1, find the variables, $\{R_{i},t_{i}\}_{i=1}^{n}\cup\{r_{ij}\}_{(i,j)\in\mathcal{E}_{r}}$ , which solve
@@
2 relaxes the special orthogonal constraint $R_{i}\in\operatorname{{\operatorname{SO}}(d)}$ to an orthogonality constraint, $R_{i}\in\operatorname{{\operatorname{O}}(d)}$, and introduces auxiliary unit-norm vector variables, $r_{ij}\in\mathbb{R}^{d}$, as a reformulation of the range cost terms. From 2, an SDP relaxation can be constructed via Shor’s relaxation [42]
@@
Find $Z\in\mathbb{R}^{k\times k}$ that solves
@@
The exact steps to form this relaxation are outlined in [9, Appendix D]. Key to our approach is recognition that 3 should often admit low-rank solutions (i.e., solutions $Z^{\ast}$ for which $\operatorname{rank}Z^{\ast}\ll k$). Indeed, this low-rank behavior has been found in practice [9]. Thus, instead of solving 3 directly via standard SDP solvers, one may instead leverage the low-rank structure by optimizing over a low-dimensional factorization $X^{\top}X=Z,X\in\mathbb{R}^{p\times k}$ for user-selectable parameter $p\ll k$. This technique, known as the Burer-Monteiro method [43, 44, 45], arrives at the following problem
@@
in which the factorization implicitly constrains the rank of the solution to be less than or equal to $p$. For our specific problem, $X$ can be structured as
@@
For higher problem dimensions ($p>d$), variables (e.g., $R_{1}$) in (5) can be thought of as higher-dimensional generalizations of the original variables of the QCQP (2); 4 thus represents a higher-dimensional relaxation of 2. In fact, analysis of the constraints of 4 reveals that the problem can be equivalently posed as Riemannian manifold optimization [9]
@@
Given the measurements of Problem 1, find the variables, $\{R_{i},t_{i}\}_{i=1}^{n}\cup\{r_{ij}\}_{(i,j)\in\mathcal{E}_{r}}$ , which solve
@@
This manifold optimization representation (5) directly generalizes the original QCQP (2), as the Stiefel manifold $\operatorname{St}(p,d)$ generalizes the orthogonal group $\operatorname{{\operatorname{O}}(d)}$ and the auxiliary variables $r_{ij}$ are still constrained to be unit-norm but can be of arbitrary dimension $p$. 5 is directly optimized in the DCORA algorithm. For notational convenience, we refer to the product manifold defined by 5’s variables as $\operatorname{\operatorname{\mathcal{M}}_{\text{RA-SLAM}}}(p,n,l)\triangleq\operatorname{St}(p,d)^{n}\times(S^{p-1})^{l}\times\mathbb{R}^{p\times n}\subset\mathbb{R}^{p\times k}$.
@@ III-B DCORA Algorithm
1:  Obtain initial point $X_{\text{init}}\in\operatorname{\operatorname{\mathcal{M}}_{\text{RA-SLAM}}}(p,n,l)$ for $p=d$ via distributed initialization.
@@
4:     Compute first-order critical point of 4 [8, Algorithm 3] $X^{\ast}\leftarrow\text{RBCD}(X)$.
@@
6:     Construct corresponding dual certificate matrix [9, Theorem 1] $S(X^{\ast})\triangleq Q+\sum_{i=1}^{m}\lambda_{i}A_{i}$, where $\lambda\in\mathbb{R}^{m}\triangleq\left[\lambda_{1},\ldots,\lambda_{m}\right]$ are the Lagrange multipliers at KKT point $X^{\ast}$.
7:     Compute minimum eigenpair [8, Algorithm 6] $(\lambda,v)\leftarrow\text{MinEig}(S(X^{\ast}))$.
@@
11:        Construct second-order descent direction $\dot{X}_{+}\triangleq[0\ v^{\top}]^{\top}$, where $v\in\mathbb{R}^{k}$ is the eigenvector corresponding to the smallest eigenvalue of $S(X^{\ast})$ which satisfies $v^{\top}S(X^{\ast})v<0$.
12:        Descend from $X^{\ast}$ [8, Algorithm 7] $X\leftarrow\text{EscapeSaddle}(X^{\ast},\dot{X}_{+})$.
@@
15:  Recover the optimal value of the SDP relaxation $f^{\ast}_{\text{SDP}}=\langle Q,{X^{\ast}}^{\top}X^{\ast}\rangle_{F}$, where $\langle\cdot,\cdot\rangle_{F}$ is the Frobenius inner product.
@@
An initial point $X_{\text{init}}\in\operatorname{\operatorname{\mathcal{M}}_{\text{RA-SLAM}}}(p=d,n,l)$ is obtained via distributed initialization. This initial point is lifted to rank $p_{0}$: $X=Y_{\text{rand}}X_{\text{init}}$, where $Y_{\text{rand}}\in\operatorname{St}(p_{0},d)$ is a random Stiefel element sampled from the initial rank $p_{0}\geq d$ of the Riemannian staircase.
From this lifted initial point, the Riemannian Staircase methodology is applied. First, 5 is optimized over via RBCD. Once RBCD has returned a first-order critical point $X^{\ast}$, $X^{\ast}$ is lifted to the next rank $p+1$ in the Riemannian staircase, where a distributed verification step either certifies $X^{\ast}$ as globally optimal or constructs a second-order descent direction to descend from $X^{\ast}$. If $X^{\ast}$ is not globally optimal, then saddle point escape is performed via line search along the computed descent direction and RBCD is used to optimize from the resulting point. The distributed Riemannian Staircase algorithm continues until a certified globally optimal value $f^{\ast}_{\text{SDP}}$ of 3 is recovered, which is a lower bound on 1’s optimal value. If the SDP is tight, as it often is in practice [9], then the solution is globally optimal for 1. Completing the DCORA algorithm, $X^{\ast}$ is projected to the feasible set of 1 via distributed rounding to recover pose and landmark estimates $T\in\operatorname{{\operatorname{SE}}(d)}^{n}\triangleq\{R_{i}\in\operatorname{{\operatorname{SO}}(d)},t_{i}\in\mathbb{R}^{d}\}_{i=1}^{n}$.
@@ III-C1 Variable Partitioning
A key notion in DCORA is that of variable ownership. Each decision variable in 5 is owned by a single agent, and variables are partitioned by agent ownership to satisfy RBCD’s input requirement. Specifically, RBCD requires a block partitioning of $X$, where a block $X_{b}\subseteq X$ is associated with the agent responsible for estimating those states. In DC2-PGO, agents naturally take ownership of pose variables that describe their trajectory. With the introduction of external landmarks and range measurements, variable partitioning becomes less obvious for the RA-SLAM problem. As such, we maintain DC2-PGO’s pose ownership rule in DCORA and assume agents own landmarks according to a consensus step [47] and auxiliary unit-norm vector variables corresponding to range measurements collected by their onboard sensors.
@@ III-C2 Variable Labeling
In an agent’s local search step, RBCD requires agents to maintain certain local copies of variables belonging to other agents. To facilitate efficient sharing of these copies, poses and landmarks are labeled public or private according to the following definition
@@ IV Experiments
Experiments conducted to evaluate DCORA are two-fold. First, we benchmark DCORA against its centralized counterpart, CORA, on public real-world multi-agent RA-SLAM datasets. Comparison is limited to CORA as (to our knowledge) this is the only other certifiably correct algorithm that can solve the general RA-SLAM problem. We refer the interested reader to [9] for an in-depth comparison of CORA against the GTSAM Levenberg–Marquardt optimizer [48], a popular local search back-end solver, to motivate our comparison. Second, multi-agent RA-SLAM datasets are generated by modifying public synthetic PGO datasets. These datasets are modified to assign pose variables to multiple agents, include landmarks, and establish inter- and intra-agent range measurement loop closures with controlled density. Parameter sweeps are performed based on (i) the number of agents, (ii) the number of landmarks, and (iii) range measurement density to study the RA-SLAM problem structure. DCORA’s implementation for evaluation precedes the discussion of these experiments.